OpenStudy (anonymous):

Ann is a plant manager in a factory. One machine on the production line breaks down frequently, stopping production. Ann kept track of the downtime of this machine for a week. Mon 75 min Tues 75 min weds 35 thurs 90 Friday 20 what is the average time the machine was down per day. I came up with 59 is this correct

OpenStudy (anonymous):

75+75=150+35=185+20=205+90=295/5= 59 you're absolutely correct :)